I’m throwing 8 dice up in the air. What’s the probability that half land on an even number, while the other half land on 1?
Answer by palanisamy(496)

8 dice are thrown.
Probability to get an even number = 3/6 = 1/2
Probability to get 1 = 1/6
Probability to get 4 even numbers and 4 one's = 8C4*(1/2)^4*(1/6)^4
= [(8*7*6*5)/(1*2*3*4)]*(1/16)*(1/1296)
= 35/10368